[figure: 9.7 cm]
MAGNETISM of Co(CO_{3})(OH_{2})_{2}(C_{3}H_{4}N_{2})_{2}

B. Świa̧tek-Tran^{a}, H. A. Kołodziej^{a}, V.H. Tran^{b}, M. Baenitz^{c}, A. Vogt^{a} aFaculty of Chemistry, University of Wrocław, 50-383 Wrocław, Poland bInstitute of Low Temperature and Structure Research, Polish Academy of Sciences, 50-950 Wrocław, Poland cMax-Planck Institut für Chemische Physik fester Stoffe, D-01187 Dresden, Germany

The carbonatodiaquobisimidazolecobalt(II) Co(CO_{3})(OH_{2})_{2}(C_{3}H_{4}N_{2})_{2} crystalizes in a monoclinic structure with the space group A2/a (Baraniak et al., J. Chem. Soc. (A) 2559 (1970)). In this contribution, we report on a detailed investigation of X-ray powder diffraction, magnetization M(T, H) and specific heat C_{p}(T, H) for a polycrystalline sample of Co(CO_{3})(OH_{2})_{2}(C_{3}H_{4}N_{2})_{2}. Based on X-ray powder diffraction data, we confirm the monoclinic structure with the following lattice parameters: a = 10.114(5), b = 7.513(3), c = 14.129(6) Å, and β = 99.564(7) ^{∘}. The magnetic susceptibility (χ = M/H) in the temperature range 20 - 400 K can be fitted with the Curie-Weiss law with the effective moment μ_{eff} = 5.2 μ_{B}/Co and the paramagnetic temperature Curie θ_{p} = - 6.9 K. Thus, the ^{4}T_{1g} ground state for Co^{2+} has found to show an unquenched spinorbit coupling and a weak antiferromagnetic interaction between Co ions. Application of magnetic field significantly affects the magnetization and specific heat at T < 20 K. Anomaly observed in C_{p}(T, H)-dependences may be interpreted in terms of a nonvibrational contribution to the thermal properties (Schottky system).

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : B. Świa̧tek-Tran

Address for correspondence : Faculty of Chemistry, University of Wrocław, 50-383 Wrocław, Poland

Email address : bstran@wchuwr.chem.uni.wroc.pl